OpenStudy (anonymous):

Use the factor theorem to prove each of the following. a. x - 3 is a factor of x3 - 9x - 9 b. x + 2 is a factor of x3 + 4x2 + 5x + 2

OpenStudy (campbell_st):

Using the factor theorem if (x - a) is a factor then f(a) = 0 so you need to find f(3) if f(x) = x^3 - 9x - 9 if f(3) = 0 then (x - 3) is a factor same process for b find f(-2) if f(-2) = 0 the (x + 2) is a factor hope it helps

OpenStudy (anonymous):

how does this prove it though? just by showing f(3) if f(x) = x3 - 9x - 9 then stating (x - 3)? do i need to show more than that?

OpenStudy (anonymous):

can someone just give me the answer and that way i can find out how it's done?

OpenStudy (campbell_st):

if there is a remainder... for f(a) then a can't be a factor... for for f(3) you have f(3) = 3^3 - 9(3) - 9 if f(3) = 0 then (x - 3) is a factor... if not its not a factor... simple as that... you could prove it by polynomial division or synthetic division but this is the simplest way

OpenStudy (campbell_st):

here is the polynomial division version  the remainder is the same as finding f(3)